### v1

Abstract
Local search is a powerful heuristic in optimization and computer science, the complexity of which has been studied in the white box and black box models. In the black box model, we are given a graph $G=(V,E)$ and oracle access to a function $f:V\to\mathbb{R}$. The local search problem is to find a vertex $v$ that is a local minimum, i.e. with $f(v)\leq f(u)$ for all $(u,v)\in E$, using as few queries to the oracle as possible. We show that if a graph $G$ admits a lazy, irreducible, and reversible Markov chain with stationary distribution $\pi$, then the randomized query complexity of local search on $G$ is $\Omega\left(\frac{\sqrt{n}}{t_{mix}\cdot\exp(3\sigma)}\right)$, where $t_{mix}$ is the mixing time of the chain and $\sigma=\max_{u,v\in V(G)}\frac{\pi(v)}{\pi(u)}.$ This theorem formally establishes a connection between the query complexity of local search and the mixing time of the fastest mixing Markov chain for the given graph. We also get several corollaries that lower bound the complexity as a function of the spectral gap, one of which slightly improves a prior result from [BCR24].
1 Introduction
Local search stands as a robust heuristic within optimization and computer science, analyzed through both white box and black box frameworks. In the black box model, we are given a graph $G=(V,E)$ alongside oracle access to a function $f:V\to R$. The objective is to identify a vertex $v$ that represents a local minimum, meaning $f(v)\leq f(u)$ for every edge $(u,v)$, while minimizing the number of vertices queried. Obtaining lower bounds for the complexity of local search has a rich history of analysis via random walks. The first pioneering work on the subject was [Ald83], which did careful tailored analysis of the hitting time of random walks on the Boolean hypercube to obtain lower bounds for local search. Another breakthrough was obtained by [Aar06], which designed a combinatorial method of obtaining lower bounds for local search inspired by the relational adversary method from quantum computing. This approach enabled obtaining sharper lower bounds for the Boolean hypercube and $d$-dimensional grid, and was successfully used in many later works. In this paper we consider the high level question: How does the geometry of the graph affect the complexity of local search? While the query complexity is comprehensively understood for neighbourhood structures such as the $d$-dimensional grid and the Boolean hypercube, knowledge remains limited for more general neighbourhood structures. Nevertheless, the spatial structure in optimization settings typically extends to more complex graphs. For instance, in scenarios such as low rank matrix estimation with data compromised by adversarial attacks, the function is defined on a Riemannian manifold rather than a traditional Euclidean space [Bon13]; thus the discretization of an optimization search space may not necessarily always correspond to some $d$-dimensional grid. For a more extensive survey on stochastic gradient descent on Riemannian manifolds, see, e.g., [Ama13]. This motivates studying local search not only on hypercubes and grids, but also on broader classes of graphs. Inspired by the observation that many lower bounds for local search are based on various types of random walks, we consider general random walks for the graph at hand and obtain lower bounds as a function of their mixing time. Our analysis uses a variant of the classical relational adversary method from [BCR24] and our main result is generic in two ways: the graph is arbitrary and the random walk evolves according to a Markov chain, which we only require to be lazy, irreducible, and reversible. This allows us to formally connect the query complexity of local search and the mixing time of the fastest mixing Markov chain for the given graph, which is a classical problem analyzed starting with [BDX04], with recent results in [OTZ21]. As a corollary, we also get a lower bound in terms of the spectral gap of the transition matrix of the chain.
2 Our Results
In this section we define the model and state our results.
2.1 Model
Let $G=(V,E)$ be a connected undirected graph and $f:V\to\mathbb{R}$ a function defined on the vertices. A vertex $v\in V$ is a local minimum if $f(v)\leq f(u)$ for all $\{u,v\}\in E$. We will write $V=[n]=\{1,\ldots,n\}$. Given as input a graph $G$ and oracle access to function $f$, the local search problem is to find a local minimum of $f$ on $G$ using as few queries as possible. Each query is of the form: “Given a vertex $v$, what is $f(v)$?”.
Query complexity.
The deterministic query complexity of a task is the total number of queries necessary and sufficient for a correct deterministic algorithm to find a solution. The randomized query complexity is the expected number of queries required to find a solution with probability at least $9/10$ for each input, where the expectation is taken over the coin tosses of the protocol.
Degree and distance.
Let $d_{max}$ and $d_{min}$ be the maximum and minimum degree of any vertex in $G$ respectively. Let $d(v)$ be the degree of $v$ for all $v\in V$. For each $u,v\in V$, let $dist(u,v)$ be the length of the shortest path from $u$ to $v$.
Markov chain.
We consider a discrete-time Markov chain on $G$ with transition matrix $\mathcal{P}$, meaning that the state space is $V$ and $\mathcal{P}_{u,v}=\mathcal{P}_{v,u}=0$ whenever $(u,v)\not\in E(G)\bigcup_{u\in V}\{\{u,u\}\}$. Suppose the chain has stationary distribution $\pi$. The chain is:
\;\;$\bullet$
lazy: if $\mathcal{P}_{u,u}\geq 1/2$ for all $u\in V$.
\;\;$\bullet$
irreducible: if all states can be reached from any starting point. Formally, for any two states $x,y\in V$ there exists an integer $t$ (possibly depending on $x$ and $y$) such that $(\mathcal{P}^{t})_{x,y}>0$, where $\mathcal{P}^{t}$ is the $t$-th power of the matrix $\mathcal{P}$.
\;\;$\bullet$
reversible: if $\pi(u)\mathcal{P}_{u,v}=\pi(v)\mathcal{P}_{v,u}$ for all $u,v\in V$.  [^4] For a formal definition of reversibility, see [LP17] equation 3.26.
For each $\epsilon>0$, the mixing time $t_{mix}(\epsilon)$ of the Markov chain [^5] For a formal definition of mixing time, see e.g. [LP17] equation 4.30. The definition in [LP17] equation 4.30 is based on the TV distance, but is equivalent to the one here by Proposition 4.2 in [LP17]. with transition matrix $\mathcal{P}$ is:
$t_{mix}(\epsilon)=\min\left\{t\in\mathbb{N}\;\Bigl{\lvert}\;\forall u\in V:\;\;\frac{1}{2}\sum_{v\in V}|(\mathcal{P}^{t})_{u,v}-\pi(v)|\leq\epsilon\right\}\,.$ (1)
2.2 Our Contributions
Our main contribution is the following theorem.
Theorem 1.
Let $G=(V,E)$ be a connected undirected graph on $n$ vertices. Consider a discrete-time, lazy, irreducible, and reversible Markov chain on $G$ with transition matrix $\mathcal{P}$ and stationary distribution $\pi$. Then the randomized query complexity of local search on $G$ is
[…]
The best lower bound given by 1 is attained by considering the Markov chain with the fastest possible mixing time for $G$; see [BDX04] for a classical reference on this problem. Indeed, for many classes of graphs there can significant gaps between the mixing time of the fastest mixing Markov chain and that of more obvious choices of Markov chains (such as the max-degree walk or the Metropolis-Hastings chain), including the barbell graph, edge-transitive graphs, and distance transitive graphs [BDPX09]. For example, when the stationary distribution is set to uniform on the barbell graph666two cliques of $n/2$ vertices connected by a single edge, the max-degree random walk has mixing time $\Theta(n^{3})$ while the fastest mixing walk mixes in only $\Theta(n^{2})$ steps (see corollaries $5.2$ and $5.3$ in [BDPX05]).
Remark 1.
Since $\sigma\geq 1$, we always have $t_{mix}(\sigma/(2n))\leq t_{mix}(1/(2n))$. Thus 1 implies the randomized query complexity is
$\Omega\left(\frac{\sqrt{n}}{t_{mix}\left(\frac{1}{2n}\right)\exp(3\sigma)}\right)\,.$
For lazy chains, the second eigenvalue is always non-negative, so Theorem 1 implies a lower bound based on the spectral gap. This leads to the following corollary:
Corollary 1.
Let $G=(V,E)$ be a connected undirected graph on $n$ vertices. Consider a discrete-time, lazy, irreducible, and reversible Markov chain on $G$ with transition matrix $\mathcal{P}$ and stationary distribution $\pi$. The randomized query complexity of local search on $G$ is
$\Omega\left(\frac{(1-\lambda_{2})\sqrt{n}}{\log(n)\exp(3\sigma)}\right),$
where $\lambda_{2}$ is the second eigenvalue of $\mathcal{P}$ and $\sigma=\max_{u,v\in V}{\pi(v)}/{\pi(u)}$.
If we constrain the ratio $d_{max}/d_{min}$ and focus on the simple lazy random walk, we can remove the dependency on $\sigma$ in 1.
Corollary 2.
Let $G=(V,E)$ be a connected undirected graph on $n$ vertices. If $d_{max}/d_{min}\leq C$ for some constant $C>0$, then the randomized query complexity of local search on $G$ is
$\Omega\left(\frac{(1-\lambda_{2})\sqrt{n}}{\log{n}}\right),$
where $\lambda_{2}$ is the second eigenvalue of the transition matrix of the simple lazy random walk on $G$.
The lower bound in 2 improves by a $\log{n}$ factor the lower bound attainable from [BCR24] for such graphs. For comparison, we state the lower bound from [BCR24] next.
Proposition 1.
[BCR24] Let $G=(V,E)$ be a connected undirected graph on $n$ vertices. If $d_{max}/d_{min}\leq C$ for some constant $C>0$, then the randomized query complexity of local search on $G$ is
$\Omega\left(\frac{(1-\lambda_{2})\sqrt{n}}{\log^{2}(n)}\right),$
where $\lambda_{2}$ is the second eigenvalue of the transition matrix of the simple lazy random walk on $G$.
The extra factor of $\log(n)$ in 1 stems from the result used to connect expansion to edge congestion (see [Chu16], corollary C.2), which rounds a fractional flow with congestion $n\log(n)$ to an integer flow with congestion $n\log^{2}(n)$. By avoiding expansion altogether, 2 also avoids this excess factor of $\log(n)$.
2.3 High Level Approach
The high level approach is as follows. We consider a set of value functions $f:V\to\mathbbm{R}$ induced by walks from a fixed starting vertex. We define $f$ such that the value at any vertex off the walk is the distance to the starting vertex of the walk, and the value at vertices on the walk is decreasing along the walk. This ensures that there is only one local minimum, namely the end of the walk. This type of construction is classical [Ald83, Aar06]. We denote the space of such functions $\mathcal{X}$. We then choose a similarity measure $r:\mathcal{X}\times\mathcal{X}\to\mathbbm{R}_{\geq 0}$ (also called the “relation”) between any two functions. Semantically, the relation measures the difficulty of distinguishing two possible input functions from each other; thus it will be useful that functions induced by walks with a long shared prefix are defined as more related. Given $r$, we invoke the relational adversary variant (Lemma 1 from [BCR24]), which implicitly defines a distribution over inputs based on $r$, and outputs a lower bound on the randomized query complexity. We choose $r$ carefully such that the distribution over walks is that of an arbitrary Markov chain. The innovation of our methodology over [BCR24] lies in the construction; where their construction is based on low-congestion paths, ours takes the more natural and general approach of using arbitrary Markov chains, including the usual lazy random walks.
Figure 1: Consider a graph $G$ with a lazy, irreducible, and reversible Markov chain $\mathcal{P}$ with stationary distribution $\pi$ and mixing time $T$. The proof fixes a walk $\mathbf{x}=[x_{0},\ldots,x_{L}]$, where $L=\lfloor\sqrt{n}\rfloor\cdot T$. The walk $\mathbf{x}$ is illustrated as a solid line, where every $T$-th node is highlighted. Sample a random walk $\mathbf{y}=[y_{0},\ldots,y_{L}]$ according to $\mathcal{P}$, conditioned on $\mathbf{y}$ and $\mathbf{x}$ having a shared prefix of length $jT$, where $j\sim U(0,\lfloor\sqrt{n}\rfloor)$. We say that $x_{jT}=y_{jT}$ is the “divergence point”. In the figure, the shared prefix of $\mathbf{x}$ and $\mathbf{y}$ is $[x_{0},\ldots,x_{T}]$, so $j=1$. A critical step of the proof is to show that no vertex $v$ is too likely to lie on $\mathbf{y}$ after its divergence from $\mathbf{x}$. To show this, we divide the walk $\mathbf{y}$ in two regions. Vertices in the region $R_{1}=[y_{jT},\ldots,y_{jT+1}]$ are collectively close to being distributed according to $\pi$. This is because the divergence point from $\mathbf{x}$ is chosen randomly. In the region $R_{2}=[y_{jT+1},...,y_{L}]$, the walk $\mathbf{y}$ has mixed, so the vertices in $R_{2}$ are close to being distributed according to $\pi$. In either case, no vertex $v$ is too likely to lie on $\mathbf{y}$ after diverging from $\mathbf{x}$.
A key step in analyzing the formula given by the relational adversary is to bound the probability that, when fixing one staircase $\mathbf{x}$ and sampling a second staircase $\mathbf{y}$ conditioned to share a uniformly randomly long initial segment with $\mathbf{x}$, no vertex $v$ is too likely to lie on the “tail” of $\mathbf{y}$ (i.e. the portion after the initial segment shared with $\mathbf{x}$). The difference between our setting and previous random walk based methods lies in this analysis. We analyze separately the portions of the tail before and after it mixes. After $\mathbf{y}$ mixes, it is of course close to uniformly random. Before $\mathbf{y}$ mixes, the randomness of the point on $\mathbf{x}$ at which $\mathbf{y}$ diverges is enough to keep $\mathbf{y}$ close enough to uniformly random. See Figure 1 for a visual depiction. This analysis is very generic, parametrized only by the mixing time of the walk used. This allows 1 to give results for walks from arbitrary Markov chains with no additional analysis (other than estimating the mixing time). The natural mixing properties of the lazy random walk on expanders then allow us to derive strong lower bounds for such graphs.
2.4 Related Work
The Boolean hypercube and the $d$-dimensional grid.
The query complexity of local search was first studied experimentally by [Tov81]. The first breakthrough in the theoretical analysis of local search was obtained by [Ald83]. Aldous stated the algorithm based on steepest descent with a warm start and showed the first nontrivial lower bound of $\Omega(2^{n/2-o(n)})$ on the query complexity for the Boolean hypercube $\{0,1\}^{n}$. The lower bound construction from [Ald83] uses Yao’s lemma and describes a hard distribution, such that if a deterministic algorithm receives a random function according to this distribution, the expected number of queries issued until finding a local minimum will be large. The random function is obtained as follows:
Consider an initial vertex $v_{0}$ uniformly at random. Set the function value at $v_{0}$ to $f(v_{0})=0$. From this vertex, start an unbiased random walk $v_{0},v_{1},\ldots$ For each vertex $v$ in the graph, set $f(v)$ equal to the first hitting time of the walk at $v$; that is, let $f(v)=\min\{t\mid v_{t}=v\}$.
The function $f$ defined this way has a unique local minimum at $v_{0}$. By a very delicate analysis of this distribution, [Ald83] showed a lower bound of $\Omega(2^{n/2-o(n)})$ on the hypercube $\{0,1\}^{n}$. This almost matches the query complexity of steepest descent with a warm start, which was also analyzed in [Ald83] and shown to take ${O}(\sqrt{n}\cdot 2^{n/2})$ queries in expectation on the hypercube. The steepest descent with a warm start algorithm applies to generic graphs too, resulting in ${O}(\sqrt{n\cdot d_{max}})$ queries overall for any graph with maximum degree $d_{max}$. Aldous’ lower bound for the hypercube was later improved by [Aar06] to $\Omega(2^{n/2}/n^{2})$ via a relational adversary method inspired from quantum computing. [Zha09] further improved this lower bound to $\Theta(2^{n/2}\cdot\sqrt{n})$ via a “clock”-based random walk construction, which avoids self-intersections. Meanwhile, [LTT89] developed a deterministic divide-and-conquer approach to solving local search that is theoretically optimal over all graphs for deterministic algorithms. On the hypercube, their method yields a lower bound of $\Omega(2^{n}/\sqrt{n})$ and an upper bound of ${O}(2^{n}\log(n)/\sqrt{n})$. For the $d$-dimensional grid $[n]^{d}$, [Aar06] used the relational adversary method there to show a randomized lower bound of $\Omega(n^{d/2-1}/\log n)$ for every constant $d\geq 3$. [Zha09] proved a randomized lower bound of $\Omega(n^{d/2})$ for every constant $d\geq 4$; this is tight as shown by Aldous’ generic upper bound. Zhang also showed improved bounds of $\Omega(n^{2/3})$ and $\Omega(n^{3/2}/\sqrt{\log n})$ for $d=2$ and $d=3$ respectively, as well as some quantum results. The work of [SY09] closed further gaps in the quantum setting as well as the randomized $d=2$ case. The problem of local search on the grid was also studied under the context of multiple limited rounds of adaptive interactions by [BL22]. [BDN19] studied the communication complexity of local search; this captures distributed settings, where data is stored on different computers.
General graphs.
More general results are few and far between. On many graphs, the simple bound from [Ald83] of $\Omega(d_{max})$ queries is the best known lower bound: hiding the local minimum in one of the $d_{max}$ leaves of a star subgraph requires checking about half the leaves in expectation to find it. [SS04] gave a quantum lower bound of $\Omega\left(\sqrt[8]{\frac{s}{d_{max}}}/\log(n)\right)$, where $s$ is the separation number of the graph. This implies the same lower bound in a randomized context, using the spectral method. Meanwhile, the best known upper bound is ${O}((s+d_{max})\cdot\log n)$ due to [SS04], which was obtained via a refinement of the divide-and-conquer procedure of [LTT89]. [DR10] studied Cayley and vertex transitive graphs and gave lower bounds for local search as a function of the number of vertices and the diameter of the graph. [Ver06] obtained upper bounds as a function of the genus of the graph. [BCR24] improved on the relational adversary method of [Aar06] and applied it to get lower bounds for local search as a function of graph congestion, with corollaries for expansion and spectral gap. We apply their methodology to an alternate construction to achieve our results, including an improvement on their spectral gap result.
Stationary points.
Local search is strongly related to the problem of local optimization where the goal is to find a stationary point of a function on $\mathbb{R}^{d}$. A standard method for solving local optimization problems is through gradient methods, which find $\epsilon$-stationary points. To show lower bounds for finding stationary points, one can similarly define a function that selects a walk in the underlying space and hides a stationary point at the end of the walk. Handling (potential) smoothness of the function and ensuring the stationary point is unique represent additional challenges. For examples of works on algorithms and complexity of computing approximate stationary points, see, e.g., [Vav93, ${}^{+}$20, CDHS20, CDHS21, BM20, DS20]).
Complexity classes.
The computational complexity of local search is captured by the class PLS, defined by [JPY88] to model the difficulty of finding locally optimal solutions to optimization problems. A related class is PPAD, introduced in [Pap94] to study the computational complexity of finding a Brouwer fixed-point. Both PLS and PPAD are subsets of the class TFNP. The class PPAD contains many natural problems that are computationally equivalent to finding a Brouwer fixed point ([CD09]), such as finding an approximate Nash equilibrium in a multiplayer or two-player game ([DGP09, CDT09]), an Arrow-Debreu equilibrium in a market ([VY11, CPY17]), and a local min-max point ([DSZ21]). The query complexity of computing an $\varepsilon$-approximate Brouwer fixed point was studied in a series of papers starting with [HPV89], which introduced a construction where the function is induced by a hidden walk. This was later improved by [CD05] and [CT07]. Recently, [FGHS22] showed that the class CLS, introduced by [DP11] to capture continuous local search, is equal to PPAD $\cap$ PLS.
3 Preliminaries
We fix a discrete-time Markov chain with transition matrix $\mathcal{P}$ that has the properties required by 1: lazy, irreducible, and reversible. Let $\pi$ denote the unique stationary distribution of the chain. For each $S\subseteq V$, let $\pi(S)=\sum_{v\in S}\pi(v)$. Moreover, let
$\sigma=\max_{u,v\in V}{\pi(v)}/{\pi(u)}\,.$ (2)
For every $k\in\mathbbm{N}$ and every walk $\mathbf{x}=(x_{0},x_{1},\ldots,x_{k})$ in $G$, let $\mathcal{P}[\mathbf{x}]$ be the probability that the random walk started at $x_{0}$ with transition matrix $\mathcal{P}$ has trajectory $\mathbf{x}$, that is:
$\mathcal{P}[\mathbf{x}]=\prod_{i=0}^{k-1}\mathcal{P}_{x_{i},x_{i+1}}\,.$ (3)
Bottleneck Ratio.
The bottleneck ratio $\Phi_{\star}$ of the Markov chain with transition matrix $\mathcal{P}$ is  [^7] See, e.g., [LP17] equation 7.5.:
$\Phi_{\star}=\min_{\begin{subarray}{c}S\subseteq V:\pi(S)\leq\frac{1}{2}\end{subarray}}\;\sum_{u\in S,v\in V\setminus S}\frac{\pi(u)\mathcal{P}_{u,v}}{\pi(S)}.$
Visiting probability.
For each pair of vertices $u,v\in V$ and integer $\ell\in\mathbbm{N}$:
  • let $\mbox{P}_{visit}(u,v,\ell)$ be the probability that a random walk that has transition matrix $\mathcal{P}$, length $\ell$, and starts at $u$ visits $v$.
  • let $\mbox{E}_{visit}(u,v,\ell)$ be the expected number of times that a random walk that has transition matrix $\mathcal{P}$, length $\ell$, and starts at $u$ visits $v$.
  • let $\mbox{P}_{end}(u,v,\ell)$ be the probability that a random walk that has transition matrix $\mathcal{P}$, length $\ell$, and starts at $u$ ends at $v$.
Edge Expansion.
Let $E(S,V\setminus S)=\{(u,v)\in E\mid u\in S,v\in V\setminus S\}$ be the set of edges with one endpoint in $S$ and the other in $V\setminus S$. The edge expansion of $G$ is
$\beta=\min_{\begin{subarray}{c}S\subseteq V:|S|\leq n/2\end{subarray}}\frac{\left|E(S,V\setminus S)\right|}{|S|}\,.$
One of the main ingredients in our proof is a variant of (classical) the relational adversary method from quantum computing given in [BCR24].
Lemma 1 ([BCR24], Theorem 3).
Consider finite sets $A$ and $B$, a set $\mathcal{X}\subseteq B^{A}$ of functions, and a map $\mathcal{H}:\mathcal{X}\to\{0,1\}$ which assigns a label to each function in $\mathcal{X}$. Additionally, we get oracle access to an unknown function $F^{*}\in\mathcal{X}$. The problem is to compute $\mathcal{H}(F^{*})$ using as few queries to $F^{*}$ as possible. [^8] In other words, we have free access to $\mathcal{H}$ and the only queries counted are the ones to $F^{*}$, which will be of the form: “What is $F^{*}(a)$?”, for some $a\in A$. The oracle will return $F^{*}(a)$ in one computational step. Let $r:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ be a non-zero symmetric function of our choice with $r(F_{1},F_{2})=0$ whenever $\mathcal{H}(F_{1})=\mathcal{H}(F_{2})$. For each $\mathcal{Z}\subseteq\mathcal{X}$, define
$M(\mathcal{Z})=\sum_{F_{1}\in\mathcal{Z}}\sum_{F_{2}\in\mathcal{X}}r(F_{1},F_{2})\;\;\;\;\text{ and }\;\;\;\;q(\mathcal{Z})=\max_{a\in A}{\sum_{F_{1}\in\mathcal{Z}}\sum_{F_{2}\in\mathcal{Z}}r(F_{1},F_{2})\cdot\mathbbm{1}_{\{F_{1}(a)\neq F_{2}(a)\}}}\,.$ (4)
If there exists a subset $\mathcal{Z}\subseteq\mathcal{X}$ with $q(\mathcal{Z})>0$, then the randomized query complexity of the problem is at least
$\min\limits_{\begin{subarray}{c}\mathcal{Z}\subseteq\mathcal{X}:q(\mathcal{Z})>0\end{subarray}}0.01\cdot{M(\mathcal{Z})}/{q(\mathcal{Z})}\,.$ (5)
4 Lower bounds via mixing time
To get lower bounds for local search, we will analyze the performance of deterministic algorithms when the input distribution is obtained by considering random functions, each of which is defined using a classical “staircase” construction. Each staircase is a random walk with transition matrix $\mathcal{P}$ on $G$. We first give the setup and then prove the main theorem.
4.1 Setup
Definition 1 (Set of walks $\mathcal{W}$ and parameter $T$).
Let $L=\lfloor\sqrt{n}\rfloor\cdot T$, where $T=t_{mix}(\frac{\sigma}{2n})$. Let $\mathcal{W}$ be the set of walks $\{\mathbf{w}\mid\mathbf{w}=(w_{0},\ldots,w_{L})\}$ in $G$ with $w_{0}$ equal to the vertex $1$ and with $\mathcal{P}_{w_{i},w_{i+1}}>0$ for all $0\leq i<L$.
Definition 2 (Milestones).
Given a walk $\mathbf{x}=(x_{0},x_{1},\ldots,x_{L})\in\mathcal{W}$, every $T$-th vertex of the walk (including the first vertex) is called a “milestone”. E.g., the first three milestones of $\mathbf{x}$ are $x_{0}$, $x_{T}$, and $x_{2T}$.
Definition 3 (Good/bad walk).
A walk $\mathbf{x}\in\mathcal{W}$ is “good” if it does not repeat any milestones and “bad” otherwise. Let $good(\mathbf{x})=True$ if $\mathbf{x}$ is good and $False$ otherwise.
Definition 4 (Heads and Tails.).
For every walk $\mathbf{x}=(x_{0},x_{1},\ldots,x_{L})\in\mathcal{W}$, let
$\begin{cases}Head(\mathbf{x},j)=(x_{0},x_{1},\ldots,x_{j\cdot T})&\forall j\in\{0,\ldots,\lfloor\sqrt{n}\rfloor\}\,.\\ Tail(\mathbf{x},j)=(x_{j\cdot T+1},x_{j\cdot T+2},\ldots,x_{L})&\forall j\in\{0,\ldots,\lfloor\sqrt{n}\rfloor\}\,.\\ Tail(\mathbf{x},j_{1},j_{2})=(x_{j_{1}\cdot T+1},x_{j_{1}\cdot T+2},\ldots,x_{j_{2}\cdot T})&\forall j_{1},j_{2}\in\{0,\ldots,\lfloor\sqrt{n}\rfloor\}\mbox{with }j_{1}\leq j_{2}\,.\end{cases}$ (6)
For all $\mathbf{x},\mathbf{y}\in\mathcal{W}$, define $J(\mathbf{x},\mathbf{y})$ as the maximum index $j$ with $Head(\mathbf{x},j)=Head(\mathbf{y},j)$.
Definition 5 (The functions $f_{\mathbf{x}}$ and $g_{\mathbf{x},b}$; the set $\mathcal{X}$).
For each walk $\mathbf{x}=(x_{0},\ldots,x_{L})\in\mathcal{W}$, define a function $f_{\mathbf{x}}:[n]\to\{-L,-L+1,\ldots,n\}$ such that for all $v\in[n]$:
$f_{\mathbf{x}}(v)=\begin{cases}dist(v,1)&\text{ if }v\not\in\mathbf{x}\\ -\max\Bigl{\{}i\in\{0,\ldots,L\}\mid x_{i}=v\Bigr{\}}&\text{ if }v\in\mathbf{x}\,.\\ \end{cases}$ (7)
For all $b\in\{0,1\}$, define $g_{\mathbf{x},b}:[n]\to\{-L,-L+1,\ldots,n\}\times\{-1,0,1\}$ so that for all $v\in[n]$:
$g_{\mathbf{x},b}(v)=\begin{cases}(f_{\mathbf{x}}(v),-1)&\text{ if }v\neq x_{L}\\ (f_{\mathbf{x}}(v),b)&\text{ if }v=x_{L}\,.\end{cases}$ (8)
Let $\mathcal{X}=\Bigl{\{}g_{\mathbf{x},b}\mid\mathbf{x}\in\mathcal{W}\text{ and }b\in\{0,1\}\Bigr{\}}$.
Definition 6 (Valid function).
Let $\mathbf{x}=(x_{0},\ldots,x_{\ell})$ be a walk in $G$. A function $f:V\to\mathbb{R}$ is valid with respect to the walk $\mathbf{x}$ if it satisfies the next conditions:
  1. 1.
    For all $u,v\in\mathbf{x}$, if $\max\{i\in\{0,\ldots,\ell\}\mid v=x_{i}\}<\max\{i\in\{0,\ldots,\ell\}\mid u=x_{i}\}$, then $f(v)>f(u)$. In other words, as one walks along the walk $\mathbf{x}$ starting from $x_{0}$ until $x_{\ell}$, if the last time the vertex $v$ appears is before the last time that vertex $u$ appears, then $f(v)>f(u)$.
  2. 2.
    For all $v\in V\setminus\mathbf{x}$, we have $f(v)=dist(x_{0},v)>0$.
  3. 3.
    $f(x_{i})\leq 0$ for all $i\in\{0,\ldots,\ell\}$.
We define a similarity measure $r$ between functions from $\mathcal{X}$, commonly referred to as the relation.
Definition 7 (The function $r$).
Let $r:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ be a symmetric function such that for each $\mathbf{x},\mathbf{y}\in\mathcal{W}$ and $b_{1},b_{2}\in\{0,1\}$,
$r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})=\begin{cases}0&\text{If }b_{1}=b_{2}\text{ or }\mathbf{x}=\mathbf{y}\text{ or }\mathbf{x}\text{ is bad or }\mathbf{y}\text{ is bad.}\\ \frac{\mathcal{P}[\mathbf{x}]\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},j)]}&\text{Otherwise, where $j=J(\mathbf{x},\mathbf{y})$}\,.\end{cases}$
Observation 1.
The function $r$ from 7 is symmetric, since by definition of $j$ we have $Head(\mathbf{x},j)=Head(\mathbf{y},j)$.
Having defined the relation $r$, applying Lemma 1 will give a lower bound by considering a distribution $p$ over $\mathcal{X}$, where each function $F\in\mathcal{X}$ is given as input with probability $p(F)=\frac{M(\{F\})}{M(\mathcal{X})},$ where $M$ is as defined in Eq. 4 for the relation $r$. What remains to be done is to explain how to invoke Lemma 1 and estimate the lower bound it gives when the input distribution is $p$.
4.2 Proof of the main theorem
See 1
Proof.
First, we may assume that $n\geq 16\sigma^{2}$. This is because in the alternative case where $n<16\sigma^{2}$, the theorem statement does not give anything useful as $\sqrt{n}<\exp(3\sigma)$ and $t_{mix}(\sigma/(2n))\geq 1$. The value functions we will use are of the form $f_{\mathbf{x}}$ as seen in Definition 5. These functions are parametrized by walks $\mathbf{x}$ of length $L$ from the set $\mathcal{W}$ defined in 1. For sake of invoking Lemma 1 however, we must turn the local search problem into a decision problem. To do this, we use the technique shown in  [Aar06, DR10]: associate with each function $f_{\mathbf{x}}$ the function $g_{\mathbf{x},b}$ defined in Definition 5. This hides a bit at the local minimum vertex (while hiding the value $-1$ at every other vertex). The decision problem is: given the graph $G$ and oracle access to a function $g_{\mathbf{x},b}$, return the hidden bit $b$; i.e. we set the function $\mathcal{H}$ for use in Lemma 1 as $\mathcal{H}(g_{\mathbf{x},b})=b$. By 2, the function $f_{\mathbf{x}}$ as defined in Definition 5 is valid. Therefore by Lemma 10, $f_{\mathbf{x}}$ has a unique local minimum, namely $x_{L}$. This means that $g_{\mathbf{x},b}$ as defined in Definition 5 does indeed hide the bit $b$ only at the local minimum of $f_{\mathbf{x}}$. Therefore measuring the query complexity of this decision problem will give the answer for local search, as the following two problems have query complexity within additive $1$:
  • search problem: given oracle access to a function $f_{x}$, find a vertex $v$ that is a local minimum;
  • decision problem: given oracle access to a function $g_{x,b}$, find $b$.
We then invoke Lemma 1 with $\mathcal{X}$ as defined in Definition 5 and with $\mathcal{H}(g_{\mathbf{x},b})=b$. This tells us that the randomized query complexity of the decision problem is at least $\min\limits_{\begin{subarray}{c}\mathcal{Z}\subseteq\mathcal{X}:q(\mathcal{Z})>0\end{subarray}}0.01\cdot{M(\mathcal{Z})}/{q(\mathcal{Z})}$, with $M(\mathcal{Z})$ and $q(\mathcal{Z})$ as defined in Lemma 1. By Lemma 7, there exists a subset $\mathcal{Z}\subseteq\mathcal{X}$ with $q(\mathcal{Z})>0$. From this point on, we fix an arbitrary subset $\mathcal{Z}\subseteq\mathcal{X}$ of functions with $q(\mathcal{Z})>0$ and will then lower bound $M(\mathcal{Z})$ and upper bound $q(\mathcal{Z})$.
Lower bounding $M(\mathcal{Z})$.
Because $q(\mathcal{Z})>0$, we know $\mathcal{Z}$ is not empty. Consider an arbitrary function $g_{\mathbf{x},b_{1}}\in\mathcal{Z}$ with $\mathbf{x}$ good. By definition of $M$, we have
$M(\{g_{\mathbf{x},b_{1}}\})$ $=\sum_{g_{\mathbf{y},b_{2}}\in\mathcal{X}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}),$ (9)
recalling:
$\bullet$
$\mathcal{P}[\mathbf{y}]=\prod_{i=0}^{k-1}\mathcal{P}_{y_{i},y_{i+1}}$ for every walk $\mathbf{y}=(y_{0},y_{1},\ldots,y_{k})$ in $G$;
$\bullet$
4 of $Head(\mathbf{y},j)$ and $J(\mathbf{x},\mathbf{y})$ and 7 of the relation $r$:
$r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})=\begin{cases}0&\text{If }b_{1}=b_{2}\text{ or }\mathbf{x}=\mathbf{y}\text{ or }\mathbf{x}\text{ is bad or }\mathbf{y}\text{ is bad.}\\ \frac{\mathcal{P}[\mathbf{x}]\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},j)]}&\text{Otherwise, where $j=J(\mathbf{x},\mathbf{y})$}\,.\end{cases}$
Then we can rewrite $M(g_{\mathbf{x},b_{1}})$ as
$M(\{g_{\mathbf{x},b_{1}}\})$ $=\sum_{\begin{subarray}{c}g_{\mathbf{y},b_{2}}\in\mathcal{X}:\\ b_{2}=1-b_{1}\\ good(\mathbf{y})\\ \mathbf{x}\neq\mathbf{y}\end{subarray}}\frac{\mathcal{P}[\mathbf{x}]\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},J(\mathbf{x},\mathbf{y}))]}\,.$ (14)
The condition $\mathbf{x}\neq\mathbf{y}$ in Eq. 14 is equivalent to $J(\mathbf{x},\mathbf{y})\neq\lfloor\sqrt{n}\rfloor$, so it must be the case that $J(\mathbf{x},\mathbf{y})\in\left\{0,\ldots,\lfloor\sqrt{n}\rfloor-1\right\}$. We decompose the summation in Eq. 14 by the value of $J(\mathbf{x},\mathbf{y})$, excluding cases where $r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})=0$, to get the following:
$M(\{g_{\mathbf{x},b_{1}}\})$ $=\mathcal{P}[\mathbf{x}]\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}g_{\mathbf{y},b_{2}}\in\mathcal{X}:\\ J(\mathbf{x},\mathbf{y})=j\\ b_{2}=1-b_{1}\\ good(\mathbf{y})\end{subarray}}\frac{\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},j)]}\,.$ (19)
There is a one-to-one correspondence between walks $\mathbf{y}\in\mathcal{W}$ and functions $g_{\mathbf{y},b_{2}}$ with $b_{2}=1-b_{1}$. Therefore we may equivalently sum over $\mathbf{y}\in\mathcal{W}$ in Eq. 19:
$M(\{g_{\mathbf{x},b_{1}}\})$ $=\mathcal{P}[\mathbf{x}]\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ J(\mathbf{x},\mathbf{y})=j\\ good(\mathbf{y})\end{subarray}}\frac{\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},j)]}\,.$ (23)
Since $\mathbf{y}$ is the concatenation of $Head(\mathbf{y},j)$ with $Tail(\mathbf{y},j)$, we have:
$\mathcal{P}[\mathbf{y}]=\mathcal{P}[Head(\mathbf{y},j)]\cdot\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}\,.$ (24)
Substituting Eq. 24 in Eq. 23 gives
$M(\{g_{\mathbf{x},b_{1}}\})$ $=\mathcal{P}[\mathbf{x}]\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ J(\mathbf{x},\mathbf{y})=j\\ good(\mathbf{y})\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}\,.$ (28)
If it weren’t for the restrictions that $J(\mathbf{x},\mathbf{y})=j$ (instead of $J(\mathbf{x},\mathbf{y})\geq j$) and $good(y)$, then the inner sum would be over all possible walks from $x_{jT}$, and would thus be $1$. Because of those restrictions, we instead invoke Lemma 5 to continue from Eq. 28 as follows:
$M(\{g_{\mathbf{x},b_{1}}\})$ $\geq\mathcal{P}[\mathbf{x}]\cdot 2^{-4\sigma}\cdot\lfloor\sqrt{n}\rfloor\,.$ (29)
If $\mathbf{x}$ is bad, then $M(\{g_{\mathbf{x},b_{1}}\})=0$, and so the function $g_{\mathbf{x},b_{1}}$ does not contribute to $M(\mathcal{Z})$. Therefore
$M(\mathcal{Z})\geq 2^{-4\sigma}\lfloor\sqrt{n}\rfloor\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]\,.$ (32)
Upper bounding $q(\mathcal{Z})$.
Define
$\widetilde{q}(\mathcal{Z},v)=\sum_{g_{\mathbf{x},b_{1}}\in\mathcal{Z}}\sum_{g_{\mathbf{y},b_{2}}\in\mathcal{Z}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})\mathbbm{1}_{\{g_{\mathbf{x},b_{1}}(v)\neq g_{\mathbf{y},b_{2}}(v)\}}\,.$ (33)
By definition of $q(\mathcal{Z})$, we have
$q(\mathcal{Z})=\max_{v\in V}\sum_{g_{\mathbf{x},b_{1}}\in\mathcal{Z}}\sum_{g_{\mathbf{y},b_{2}}\in\mathcal{Z}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})\mathbbm{1}_{\{g_{\mathbf{x},b_{1}}(v)\neq g_{\mathbf{y},b_{2}}(v)\}}=\max_{v\in V}\widetilde{q}(\mathcal{Z},v)\,.$ (34)
We will bound $q(\mathcal{Z})$ by bounding $\widetilde{q}(\mathcal{Z},v)$ for an arbitrary $v$. Fix an arbitrary vertex $v$. We first partition the inner sum according to $J(\mathbf{x},\mathbf{y})$:
$\widetilde{q}(\mathcal{Z},v)$ $=\sum_{g_{\mathbf{x},b_{1}}\in\mathcal{Z}}\sum_{g_{\mathbf{y},b_{2}}\in\mathcal{Z}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})\mathbbm{1}_{\{g_{\mathbf{x},b_{1}}(v)\neq g_{\mathbf{y},b_{2}}(v)\}}$ (By definition of $\widetilde{q}(\mathcal{Z},v)$.)
$=\sum_{g_{\mathbf{x},b_{1}}\in\mathcal{Z}}\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ J(\mathbf{x},\mathbf{y})=j\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})\mathbbm{1}_{\{g_{\mathbf{x},b_{1}}(v)\neq g_{\mathbf{y},b_{2}}(v)\}}\,.$ (37)
By Lemma 6, we may continue from Eq. 37 and expand to get
$\widetilde{q}(\mathcal{Z},v)$ $\leq 2\sum_{g_{\mathbf{x},b_{1}}\in\mathcal{Z}}\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{y},j)\\ J(\mathbf{x},\mathbf{y})=j\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})$ (41)
$=2\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{y},j)\\ J(\mathbf{x},\mathbf{y})=j\\ b_{2}=1-b_{1}\\ good(\mathbf{y})\end{subarray}}\frac{\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},j)]}$ (Using the definition of $r$.)
$\leq 2\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}g_{\mathbf{y},b_{2}}\in\mathcal{X}:\\ v\in Tail(\mathbf{y},j)\\ J(\mathbf{x},\mathbf{y})=j\\ b_{2}=1-b_{1}\\ good(\mathbf{y})\end{subarray}}\frac{\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},j)]}\,.$ (Since $\mathcal{Z}\subseteq\mathcal{X}$.)
Again, there is a one-to-one correspondence between walks $\mathbf{y}\in\mathcal{W}$ and functions $g_{\mathbf{y},b_{2}}\in\mathcal{X}$ with $b_{2}=1-b_{1}$, so we may equivalently sum over $\mathbf{y}\in\mathcal{W}$ in LABEL:eq:q_analysis_2_2. Additionally, we expand the scope from $J(\mathbf{x},\mathbf{y})=j$ to $J(\mathbf{x},\mathbf{y})\geq j$, which is equivalent to $Head(\mathbf{x},j)=Head(\mathbf{y},j)$. Finally, we drop the requirement that $good(\mathcal{Y})$. Continuing from LABEL:eq:q_analysis_2_2,
$\widetilde{q}(\mathcal{Z},v)$ $\leq 2\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\frac{\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},j)]}\,.$ (62)
From here we partition based on where in $Tail(\mathbf{y},j)$ the vertex $v$ lies: the first short part of the tail, i.e. $Tail(\mathbf{y},j,j+1)$, or the rest of the tail, i.e. $Tail(\mathbf{y},j+1)$. Continuing from Eq. 62,
$\widetilde{q}(\mathcal{Z},v)$ $\leq 2\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\left(\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j,j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\frac{\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},j)]}+\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\frac{\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},j)]}\right)$ (71)
Since $\mathbf{y}$ is the concatenation of $Head(\mathbf{y},j)$ with $Tail(\mathbf{y},j)$, we have
$\mathcal{P}[\mathbf{y}]=\mathcal{P}[Head(\mathbf{y},j)]\cdot\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}\,.$ (72)
Using Eq. 72 in Eq. 71, we obtain
$\begin{split}\widetilde{q}(\mathcal{Z},v)\leq 2\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}&\Biggl{(}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j,j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}+\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}\Biggr{)}\end{split}$ (73)
To bound the first part, we will use Lemma 4. Since $\mathbf{x}$ is good, we have
$\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j,j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}$ $=\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\mbox{P}_{visit}(x_{jT},v,T)$ (By definition of $\mbox{P}_{visit}$.)
$\leq T\sigma\,.$ (By Lemma 4 since $\mathbf{x}$ is good.)
To bound the second part, we first partition according to the possible values of $\mathbf{y}_{(j+1)T}$. We have
$\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}$ $=\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{u\in V}\mbox{P}_{end}(x_{jT},u,T)\mbox{P}_{visit}(u,v,L-(j+1)T)$ (81)
The visiting probability $\mbox{P}_{visit}(u,v,\ell)$ is increasing in $\ell$, so continuing from Eq. 81 we get
$\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}$ (85)
$\qquad\leq\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{u\in V}\mbox{P}_{end}(x_{jT},u,T)\mbox{P}_{visit}(u,v,L)$ (86)
$\qquad\leq\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{u\in V}\Bigl{(}\pi(u)+|\mbox{P}_{end}(x_{jT},u,T)-\pi(u)|\Bigr{)}\mbox{P}_{visit}(u,v,L),$ (87)
where in Eq. 87 we used the inequality $a\leq b+|a-b|$ for $a,b\geq 0$. A random walk with starting vertex drawn from $\pi$ has probability $\pi(v)$ of being at $v$ at each step. Formally, for all $\ell\in\mathbbm{N}$ we have
$\sum_{u\in V}\pi(u)\mbox{P}_{end}(u,v,\ell)=\pi(v)\,.$ (88)
Additionally, by union bound we have
$\mbox{P}_{visit}(u,v,L)\leq\sum_{\ell=1}^{L}\mbox{P}_{end}(u,v,\ell)\,.$ (89)
We have
$\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{u\in V}\left(\pi(u)+|\mbox{P}_{end}(x_{jT},u,T)-\pi(u)|\right)\mbox{P}_{visit}(u,v,L)$ (90)
$\qquad\leq\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\left(\sum_{\ell=1}^{L}\sum_{u\in V}\pi(u)\mbox{P}_{end}(u,v,\ell)+\sum_{u\in V}|\mbox{P}_{end}(x_{jT},u,T)-\pi(u)|\cdot\mbox{P}_{visit}(u,v,L)\right)$ (By Eq. 89.)
$\qquad=\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\left(L\pi(v)+\sum_{u\in V}|\mbox{P}_{end}(x_{jT},u,T)-\pi(u)|\cdot\mbox{P}_{visit}(u,v,L)\right)\,.$ (By Eq. 88)
Since $T=t_{mix}(\sigma/(2n))$, we have
$\sum_{u\in V}|\mbox{P}_{end}(x_{jT},u,T)-\pi(u)|\cdot\mbox{P}_{visit}(u,v,L)$ $\leq\max_{u\in V}\mbox{P}_{visit}(u,v,L)\sum_{u\in V}|\mbox{P}_{end}(x_{jT},u,T)-\pi(u)|$
$=\max_{u\in V}\mbox{P}_{visit}(u,v,L)\sum_{u\in V}|(\mathcal{P}^{T})_{x_{jT},u}-\pi(u)|$
$\leq\max_{u\in V}\mbox{P}_{visit}(u,v,L)\cdot\frac{\sigma}{n}\,.$ (92)
Combining LABEL:eq:474673373 and Eq. 92 yields
$\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{u\in V}\left(\pi(u)+|\mbox{P}_{end}(x_{jT},u,T)-\pi(u)|\right)\mbox{P}_{visit}(u,v,L)\leq\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\left(L\pi(v)+\frac{\sigma}{n}\max_{u\in V}\mbox{P}_{visit}(u,v,L)\right)$ (93)
$\qquad\leq\lfloor\sqrt{n}\rfloor(L+1)\frac{\sigma}{n}$ (Since $\mbox{P}_{visit}(u,v,L)\leq 1$ and $\pi(v)\leq\frac{\sigma}{n}$.)
$\qquad\leq 2T\sigma\,.$ (Since $L=T\lfloor\sqrt{n}\rfloor$ and $1\leq L$.)
Combining Eq. 87 and LABEL:eq:second_segment, we obtain
$\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}\leq 2T\sigma\,.$ (98)
Combining Eq. 73, LABEL:eq:first_segment, and Eq. 98,
$\widetilde{q}(\mathcal{Z},v)$ $\leq 2\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]3T\sigma=6T\sigma\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]\,.$ (103)
Since $q(\mathcal{Z})=\min_{v\in V}\widetilde{q}(\mathcal{Z},v)$ and Eq. 103 holds for an arbitrary choice of $v$, we get
$q(\mathcal{Z})\leq 6T\sigma\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]\,.$ (106)
Bounding ${M(\mathcal{Z})}/{q(\mathcal{Z})}$.
Combining Eq. 32 and Eq. 106, we obtain
$\frac{M(\mathcal{Z})}{q(\mathcal{Z})}\geq\frac{2^{-4\sigma}\lfloor\sqrt{n}\rfloor\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]}{6T\sigma\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}}\in\mathcal{Z}\\ good(\mathbf{x})\end{subarray}}\mathcal{P}[\mathbf{x}]}=\frac{2^{-4\sigma}}{6\sigma}\cdot\frac{\lfloor\sqrt{n}\rfloor}{T}\,.$ (111)
We have $2^{-4\sigma}/(6\sigma)\in\Omega(1/exp(3\sigma))$. Thus applying Lemma 1 to the expression in Eq. 111, we obtain the required lower bound for local search on $G$, namely
$\Omega\left(\frac{\sqrt{n}}{t_{mix}\left(\frac{\sigma}{2n}\right)\cdot\exp(3\sigma)}\right)\,.$
This completes the proof. ∎
Appendix A Helper lemmas
Lemma 2.
For each walk $\mathbf{x}=(x_{0},\ldots,x_{L})\in\mathcal{W}$, the function $f_{\mathbf{x}}$ given by Definition 5 is valid for the walk ${\mathbf{x}}$.
Proof.
We show the conditions required by the definition of a valid function (Definition 6) hold.
Condition 1.
Consider two arbitrary vertices $v_{1},v_{2}\in{\mathbf{x}}$. Define
$i_{1}=\max\bigl{\{}k\in\{0,\ldots,L\}\mid v_{1}=x_{k}\bigr{\}}\qquad\mbox{and}\qquad i_{2}=\max\bigl{\{}k\in\{0,\ldots,L\}\mid v_{2}=x_{k}\bigr{\}}\,.$
Without loss of generality, we have $i_{1}\leq i_{2}$. Condition $1$ requires that if $i_{1}<i_{2}$, then $f_{\mathbf{x}}(v_{1})>f_{\mathbf{x}}(v_{2})$. Since $v_{1},v_{2}\in\mathbf{x}$, Definition 5 of the function $f_{\mathbf{x}}$ states that $f_{\mathbf{x}}(v_{1})=-i_{1}$ and $f_{\mathbf{x}}(v_{2})=-i_{2}$. Thus if $i_{1}<i_{2}$, then $f_{\mathbf{x}}(v_{1})>f_{\mathbf{x}}(v_{2})$, as required.
Condition 2.
The second condition requires that $f_{\mathbf{x}}(v)=dist(x_{0},v)>0$ for all $v\in V\setminus\mathbf{x}$. Since $\mathbf{x}\in\mathcal{W}$, we have $x_{0}=1$. Using Definition 5 we have $f_{\mathbf{x}}(v)=dist(v,1)=dist(v,x_{0})>0$ for all $v\not\in\mathbf{x}$.
Condition 3.
The third condition requires that $f_{\mathbf{x}}(x_{i})\leq 0$ for all $i\in\{0,\ldots,L\}$. This condition is clearly met since $f_{\mathbf{x}}(v)\in\{0,-1,\ldots,-L\}$ for all $v\in\mathbf{x}$.
Therefore $f_{\mathbf{x}}$ is valid for the walk ${\mathbf{x}}$. ∎
Lemma 3.
Let $(w_{0},w_{1},\ldots)$ be a Markov chain generated by transition matrix $\mathcal{P}$ with arbitrary starting distribution. Then for all $v\in V$ we have
$\Pr[w_{t_{mix}(\sigma/(2n))}=v]\leq 2\sigma/n\,.$
Proof.
By definition of $t_{mix}(\sigma/(2n))$ we have
$\sum_{v\in V}\left|\Pr[w_{t_{mix}(\sigma/(2n))}=v]-\pi(v)\right|\leq\sigma/n\,.$ (112)
Since each term of the sum in Eq. 112 is non-negative, we have
$\forall v\in V\;\left|\Pr[w_{t_{mix}(\sigma/(2n))}=v]-\pi(v)\right|\leq\sigma/n\,.$ (113)
Removing the absolute value from Eq. 113 and rearranging, we get
$\forall v\in V\;\Pr[w_{t_{mix}(\sigma/(2n))}=v]$ $\leq\sigma/n+\pi(v)$
$\leq 2\sigma/n\,.$ (Since $\sigma\geq\pi(v)\cdot n$.)
This concludes the proof of the lemma. ∎
Lemma 4.
Let $S\subseteq V$ be a subset of vertices. Let $v\in V$ be a vertex and $\ell\in\mathbbm{N}$. Then
$\sum_{u\in S}\mbox{\em P}_{visit}(u,v,\ell)\leq\ell\sigma\,.$
Proof.
Let $T_{\ell}$ be the random variable representing the number of times a random walk of length $\ell$ starting at $v$ visits a vertex in $S$. Decomposing by vertices in $S$, we have
$\sum_{u\in S}\mbox{P}_{visit}(u,v,\ell)$ $\leq\sum_{u\in S}\mbox{E}_{visit}(u,v,\ell)$ (114)
$=\sum_{u\in S}\mbox{E}_{visit}(v,u,\ell)\frac{\pi(v)}{\pi(u)}\,.$ (By Lemma 9)
Using the definition of $\sigma=\max_{u,v\in V}{\pi(v)}/{\pi(u)}$, we have
$\sum_{u\in S}\mbox{E}_{visit}(v,u,\ell)\frac{\pi(v)}{\pi(u)}$ $\leq\sigma\sum_{u\in S}\mbox{E}_{visit}(v,u,\ell)$
$={\sigma\cdot\operatorname{\mathbb{E}}[T_{\ell}]}$ (By definition of $T_{\ell}$.)
$\leq\ell\sigma\,.$ (Since $T_{\ell}\leq\ell$.)
Combining LABEL:eq:bounding_sum_of_p_visit_by_expected_visit_times_ratio_fractions and LABEL:eq:bounding_sum_of_expected_visits_over_S, we get $\sum_{u\in S}\mbox{P}_{visit}(u,v,\ell)\leq\ell\sigma$, as required. ∎
Lemma 5.
Let $n\geq 16\sigma^{2}$. Fix a good walk $\mathbf{x}=(x_{0},\ldots,x_{L})$ with $x_{0}=1$ and $\mathcal{P}_{x_{i},x_{i+1}}>0$ for all $0\leq i<L$. Then for each $0\leq j<\lfloor\sqrt{n}\rfloor$,
$\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ J(\mathbf{x},\mathbf{y})=j\\ good(\mathbf{y})\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}\geq 2^{-4\sigma}\,.$ (120)
Proof.
Let $P_{\mathcal{W}}$ be the distribution over the set of walks $\mathcal{W}$ generated by sampling a walk according to $\mathcal{P}$ starting at the vertex $1$ with $L=\lfloor\sqrt{n}\rfloor\cdot T$ edges. Let $\mathbf{z}$ be a random walk drawn from $P_{\mathcal{W}}$. Recall that every $T$-th vertex of a walk is called a milestone, and a walk is good if it does not repeat milestones. We have
$\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ J(\mathbf{x},\mathbf{y})=j\\ good(\mathbf{y})\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}$ $=\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ J(\mathbf{x},\mathbf{y})=j\\ good(\mathbf{y})\end{subarray}}\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}\cdot\mathcal{P}_{y_{j\cdot T+1},y_{j\cdot T+2}}\cdot\ldots\cdot\mathcal{P}_{y_{L-1},y_{L}}$ (127)
$=\Pr\Bigl{[}good(\mathbf{z})\wedge J(\mathbf{x},\mathbf{z})=j\;|\;Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}\,.$ (128)
Figure 2: The milestone $z_{(j+1)T}$, shown in purple, may not match any of the orange milestones: $x_{0}$ through $x_{jT}$ because it would make $\mathbf{z}$ bad, and $x_{(j+1)T}$ because it would make $J(\mathbf{x},\mathbf{z})>j$.
Figure 2: The milestone $z_{(j+1)T}$, shown in purple, may not match any of the orange milestones: $x_{0}$ through $x_{jT}$ because it would make $\mathbf{z}$ bad, and $x_{(j+1)T}$ because it would make $J(\mathbf{x},\mathbf{z})>j$.
Figure 3: The milestone $z_{(j+3)T}$, shown in purple, may not match any of the orange milestones because it would make $\mathbf{z}$ bad.
Equivalently, we can sample $\mathbf{z}\sim P_{\mathcal{W}}$ one segment at a time with the constraint that the initial $jT+1$ vertices of $\mathbf{z}$ must match those of $\mathbf{x}$. That is, set $Head(\mathbf{z},j)=Head(\mathbf{x},j)$. Then for
$i=j,\ldots,\lfloor\sqrt{n}\rfloor$, sample the segment $Tail(\mathbf{z},i-1,i)$ conditioned on having set $Head(\mathbf{z},i-1)$. For every $0\leq k\leq\lfloor\sqrt{n}\rfloor$, the set of vertices given by the first $k+1$ milestones of $\mathbf{z}$ is
$S_{k}=\{z_{0},z_{T},\ldots,z_{kT}\}\,.$ (129)
Given that $Head(\mathbf{x},j)=Head(\mathbf{z},j)$, the condition $J(\mathbf{x},\mathbf{z})=j$ is equivalent to
$z_{(j+1)T}\neq x_{(j+1)T}\,.$ (130)
Similarly, given that $Head(\mathbf{x},j)=Head(\mathbf{z},j)$, the condition that $\mathbf{z}$ is good is equivalent to
$z_{kT}\not\in S_{k-1}\qquad\forall k\in\{j+1,j+2,\ldots,\lfloor\sqrt{n}\rfloor\}\,.$ (131)
For each $i\in\{j+1,\ldots,\lfloor\sqrt{n}\rfloor$, define
$Q_{i}=\begin{cases}\{x_{(j+1)T}\}\cup S_{j}&\text{if}\;\;i=j+1\\ S_{i-1}&\text{if}\;\;j+1<i\leq\lfloor\sqrt{n}\rfloor\,.\end{cases}$ (132)
Combining (130) and (131), we get
$\Pr\Bigl{[}good(\mathbf{z})\wedge J(\mathbf{x},\mathbf{z})=j\;|\;Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}$
$\qquad=\Pr\Bigl{[}\Bigl{(}z_{kT}\not\in S_{k-1}\;\forall k\in\{j+1,\ldots,\lfloor\sqrt{n}\rfloor\}\Bigr{)}\land\Bigl{(}z_{(j+1)T}\neq x_{(j+1)T}\ \Bigr{)}\mid Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}$ (133)
$\qquad=\Pr\left[\bigwedge_{i=j+1}^{\lfloor\sqrt{n}\rfloor}z_{iT}\notin Q_{i}\mid Head(\mathbf{x},j)=Head(\mathbf{z},j)\right]$ (By definition of $Q_{i}$ in (132))
$\qquad=\prod_{i=j+1}^{\lfloor\sqrt{n}\rfloor}\Pr\left[z_{iT}\notin Q_{i}\mid\left(Head(\mathbf{x},j)=Head(\mathbf{z},j)\right)\wedge\bigl{(}\bigwedge_{k=j+1}^{i-1}z_{kT}\notin Q_{k}\bigr{)}\right]$ (134)
For all $i\leq\lfloor\sqrt{n}\rfloor$, let $\mathcal{W}_{i}$ be the space of walks of length $Ti$ that can occur with positive probability under transition matrix $\mathcal{P}$, formally defined as:
$\mathcal{W}_{i}=\Bigl{\{}\mathbf{w}\mid\mathbf{w}=(w_{0},\ldots,w_{T\cdot i})\text{ where }w_{0}=1\text{ and }\mathcal{P}_{w_{k},w_{k+1}}>0\text{ for all }0\leq k<T\cdot i\Bigr{\}}\,.$ (135)
Then, using Appendix A gives
$\Pr\Bigl{[}good(\mathbf{z})\wedge J(\mathbf{x},\mathbf{z})=j\;|\;Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}$
$\qquad\geq\prod_{i=j+1}^{\lfloor\sqrt{n}\rfloor}\min_{\mathbf{\xi}\in\mathcal{W}_{i-1}}\Pr\Bigl{[}z_{iT}\notin Q_{i}\mid Head(\mathbf{z},i-1)=\mathbf{\xi}\Bigr{]}\,.$ (136)
Since $T=t_{mix}(\sigma/(2n))$, Lemma 3 tells us that for all $v\in V$ and $1\leq i\leq\lfloor\sqrt{n}\rfloor$
$\min_{\xi\in\mathcal{W}_{i-1}}\Pr\Bigl{[}z_{iT}=v\mid Head(\mathbf{z},i-1)=\xi\Bigr{]}\leq\frac{2\sigma}{n}\,.$ (137)
By the union bound applied to Eq. 137, for each set $R\subseteq V$ and $1\leq i\leq\lfloor\sqrt{n}\rfloor$, we have
$\min_{\xi\in\mathcal{W}_{i-1}}\Pr\Bigl{[}z_{iT}\notin R\mid Head(\mathbf{z},i-1)=\xi\Bigr{]}\geq 1-\frac{2\sigma|R|}{n}\,.$ (138)
We consider two cases based on $j$:
Case 1: $j<\lfloor\sqrt{n}\rfloor-1$.
Since $j<\lfloor\sqrt{n}\rfloor-1$, we have
|$|\{x_{(j+1)T}\}\cup S_{j}|\leq\lfloor\sqrt{n}\rfloor\,.$ (139)
Furthermore, for all $j<k<\lfloor\sqrt{n}\rfloor$ we have
$|S_{k}|\leq k+1\leq\lfloor\sqrt{n}\rfloor\,.$ (140)
Combining Eq. 139 and Eq. 140 gives
$|Q_{i}|\leq\lfloor\sqrt{n}\rfloor\qquad\forall j+1\leq i\leq\lfloor\sqrt{n}\rfloor\,.$ (141)
We obtain:
$\Pr\Bigl{[}good(\mathbf{z})\wedge J(\mathbf{x},\mathbf{z})=j\;|\;Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}$
$\qquad\geq\prod_{i=j+1}^{\lfloor\sqrt{n}\rfloor}\min_{\mathbf{\xi}\in\mathcal{W}_{i-1}}\Pr\Bigl{[}z_{iT}\notin Q_{i}\mid Head(\mathbf{z},i-1)=\mathbf{\xi}\Bigr{]}$ (By Appendix A)
$\qquad\geq\left(1-\frac{2\sigma\lfloor\sqrt{n}\rfloor}{n}\right)^{\lfloor\sqrt{n}\rfloor-j}$ (By Eq. 138 with $R=Q_{i}$ and since $|Q_{i}|\leq\sqrt{n}$ by Eq. 141.)
$\qquad\geq\left(1-\frac{2\sigma}{\sqrt{n}}\right)^{\sqrt{n}}\,.$ (142)
By Lemma 8, we have that $\left(1-\frac{2\sigma}{\sqrt{n}}\right)^{\sqrt{n}}$ is an increasing function of $\sqrt{n}$ since $n\geq 16\sigma^{2}$. Therefore it is minimized at $\sqrt{n}=4\sigma$. Substituting in Eq. 142, we get
$\Pr\Bigl{[}good(\mathbf{z})\wedge J(\mathbf{x},\mathbf{z})=j\;|\;Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}$ $\geq\left(1-\frac{2\sigma}{4\sigma}\right)^{4\sigma}=2^{-4\sigma}\,.$ (143)
Case 2: $j=\lfloor\sqrt{n}\rfloor-1$.
We again invoke Eq. 138. In this case,
$|Q_{\lfloor\sqrt{n}\rfloor}|\leq|S_{\lfloor\sqrt{n}\rfloor-1}|+1\leq\sqrt{n}+1\,.$ (144)
Using Appendix A, we obtain:
$\Pr\Bigl{[}good(\mathbf{z})\wedge J(\mathbf{x},\mathbf{z})=j\;|\;Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}$
$\qquad\geq\min_{\mathbf{\xi}\in\mathcal{W}_{\lfloor\sqrt{n}\rfloor-1}}\Pr\Bigl{[}z_{\lfloor\sqrt{n}\rfloor T}\notin Q_{\lfloor\sqrt{n}\rfloor}\mid Head(\mathbf{z},\lfloor\sqrt{n}\rfloor-1)=\mathbf{\xi}\Bigr{]}\,.$ (145)
Using Eq. 138 with $R=Q_{\lfloor\sqrt{n}\rfloor}$ and since $|Q_{\lfloor\sqrt{n}\rfloor}\leq\sqrt{n}+1$ by Eq. 144, we can lower bound the right hand side of Eq. 145 as follows:
$\min_{\mathbf{\xi}\in\mathcal{W}_{\lfloor\sqrt{n}\rfloor-1}}\Pr\Bigl{[}z_{\lfloor\sqrt{n}\rfloor T}\notin Q_{\lfloor\sqrt{n}\rfloor}\mid Head(\mathbf{z},\lfloor\sqrt{n}\rfloor-1)=\mathbf{\xi}\Bigr{]}$ $\geq 1-\frac{2\sigma(\lfloor\sqrt{n}\rfloor+1)}{n}\,.$ (146)
Combining Eq. 145 and Eq. 146, we get
$\Pr\Bigl{[}good(\mathbf{z})\wedge J(\mathbf{x},\mathbf{z})=j\;|\;Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}\geq 1-\frac{2\sigma(\lfloor\sqrt{n}\rfloor+1)}{n}\,.$ (147)
We can further lower bound the right hand side of Eq. 147 as follows:
$1-\frac{2\sigma(\lfloor\sqrt{n}\rfloor+1)}{n}$ $\geq 1-\frac{\sqrt{n}(\sqrt{n}+1)}{2n}$ (Since $\sqrt{n}\geq 4\sigma$.)
$\geq\frac{1}{4}$ (Since $n\geq 16\sigma^{2}\geq 16$.)
$\geq 2^{-4\sigma}\,.$ (Since $\sigma\geq 1$.)
Combining Eq. 147 and Eq. 149, we get
$\Pr\Bigl{[}good(\mathbf{z})\wedge J(\mathbf{x},\mathbf{z})=j\;|\;Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}\geq 2^{-4\sigma}\,.$ (149)
In both cases, we obtained (by Eq. 143 and Eq. 149)
$\Pr\Bigl{[}good(\mathbf{z})\wedge J(\mathbf{x},\mathbf{z})=j\;|\;Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}\geq 2^{-4\sigma}\,.$ (150)
Combining Eq. 127 with Eq. 150 yields
$\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ J(\mathbf{x},\mathbf{y})=j\\ good(\mathbf{y})\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}\geq 2^{-4\sigma}\,.$ (154)
This concludes the proof of the lemma. ∎ The next lemma is inspired by Lemma 9 from [BCR24]. However, it is slightly different, so we include the proof here.
Lemma 6.
Let $v\in V$ and $\mathcal{Z}\subseteq\mathcal{X}$. Then we have
$\sum_{g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})\mathbbm{1}_{\{g_{\mathbf{x},b_{1}}(v)\neq g_{\mathbf{y},b_{2}}(v)\}}\leq 2\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{y},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})\,.$
Proof.
If $g_{\mathbf{x},b_{1}}(v)\neq g_{\mathbf{y},b_{2}}(v)$, then either:
  • $v\in Tail(\mathbf{x},J(\mathbf{x},\mathbf{y}))\cup Tail(\mathbf{y},J(\mathbf{x},\mathbf{y}))$
  • or $\mathbf{x}=\mathbf{y}$, in which case $r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})=0$.
Therefore
$\sum_{g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})\mathbbm{1}_{\{g_{\mathbf{x},b_{1}}(v)\neq g_{\mathbf{y},b_{2}}(v)\}}$ $\leq\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{x},J(\mathbf{x},\mathbf{y}))\cup Tail(\mathbf{y},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})$ (157)
$\leq\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{x},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})+\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{y},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})$ (162)
$=2\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{y},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})$ (By symmetry of $r$.)
This completes the proof of the lemma. ∎
Lemma 7.
In the setting of Theorem 1, if $n\geq 16\sigma^{2}$ then there exists a subset $\mathcal{Z}\subseteq\mathcal{X}$ with $q(\mathcal{Z})>0$.
Proof.
By definition of $\sigma$ and because $\max_{v\in V}\pi(v)\geq 1/n$, we have
$\min_{v}\pi(v)\geq 4/n^{1.5}\,.$ (165)
Fix an arbitrary vertex $u$. Let $S$ be the set of vertices unreachable from $u$ via a random walk that evolves according to $\mathcal{P}$ and has at most $T=t_{mix}(\frac{\sigma}{2n})$ steps. Denote $s=|S|$. Then
$(\mathcal{P}^{T})_{u,v}=0\;\;\forall v\in S\,.$ (166)
By definition of $t_{mix}$, we have
$\frac{\sigma}{2n}$ $\geq\frac{1}{2}\sum_{v\in V}|(\mathcal{P}^{T})_{u,v}-\pi(v)|$
$\geq\frac{1}{2}\sum_{v\in S}|\pi(v)|$ (By Eq. 166)
$\geq\frac{s}{2}\min_{v\in V}\pi(v)$ (Since $s=|S|$.)
$\geq\frac{2s}{n^{1.5}}\,.$ (By Eq. 165.)
Meanwhile, since $n\geq 16\sigma^{2}$, we have
$\frac{\sigma}{2n}\leq\frac{1}{8\sqrt{n}}\,.$ (168)
Combining LABEL:eq:exists_Z_1 and Eq. 168, we get $s\leq n/16$. Thus the number of vertices reachable from $u$ in $T$ steps is $n-s\geq 15n/16$. For $n\geq 5$, we have $215n/16\geq\lfloor\sqrt{n}\rfloor+$ 2, which means that at least $\lfloor\sqrt{n}\rfloor+2$ vertices are reachable via $\mathcal{P}$ from any vertex $u$ within $T$ steps. We can then construct two walks, $\mathbf{x}$ and $\mathbf{y}$, in the following manner. For $i=1,\ldots,\lfloor\sqrt{n}\rfloor$, take $x_{iT}$ to be an arbitrary vertex reachable from $x_{(i-1)T}$ other than $x_{0},\ldots,x_{(i-1)T}$; this is possible since at least $\lfloor\sqrt{n}\rfloor+2$ vertices are reachable. Connect the milestones using an arbitrary path such that every edge $(u,w)$ in the path has $\mathcal{P}_{u,w}>0$. Construct $\mathbf{y}$ in the same manner but requiring that $Head(\mathbf{y},\lfloor\sqrt{n}\rfloor-1)=Head(\mathbf{x},\lfloor\sqrt{n}\rfloor-1)$ and $y_{L}\not\in\left\{x_{0},x_{T},x_{2T},\ldots,x_{\lfloor\sqrt{n}\rfloor T}\right\}$. Define $\mathcal{Z}=\{g_{\mathbf{x},0},g_{\mathbf{y},1}\}$. We will show that $q(\mathcal{Z})>0$. First observe that $\mathbf{x}\neq\mathbf{y}$, both $\mathbf{x}$ and $\mathbf{y}$ are good since they do not repeat milestones, and the bit hidden by $g_{\mathbf{x},0}$ is different from the bit hidden by $g_{\mathbf{y},1}$. The length of the prefix shared by $\mathbf{x}$ and $\mathbf{y}$ is $J(\mathbf{x},\mathbf{y})=\lfloor\sqrt{n}\rfloor-1$. Then
$r(g_{\mathbf{x},0},g_{\mathbf{y},1})=\frac{\mathcal{P}[\mathbf{x}]\mathcal{P}[\mathbf{y}]}{\mathcal{P}[Head(\mathbf{y},\lfloor\sqrt{n}\rfloor-1)]}>0\,.$ (169)
We have
$q(\mathcal{Z})$ $=\max_{v\in V}\sum_{F_{1}\in\mathcal{Z}}\sum_{F_{2}\in\mathcal{Z}}r(F_{1},F_{2})\cdot\mathbbm{1}_{\{F_{1}(v)\neq F_{2}(v)\}}$
$=2\max_{v\in V}r(g_{\mathbf{x},0},g_{\mathbf{y},1})\cdot\mathbbm{1}_{\{g_{\mathbf{x},0}(v)\neq g_{\mathbf{y},1}(v)\}}$ (Since $\mathcal{Z}=\{g_{\mathbf{x},0},g_{\mathbf{y},1}\}$ and $r(F,F)=0$ $\forall F\in\mathcal{Z}$.)
$>0\,.$ (Using Eq. 169 and $g_{\mathbf{x},0}(x_{L})\neq g_{\mathbf{y},1}(y_{L})$.)
This completes the proof. ∎
Lemma 8.
For all $x\geq 2y\geq 1$ we have
$\frac{\partial}{\partial x}\left(1-\frac{y}{x}\right)^{x}\geq 0\,.$ (170)
Proof.
Define $z=x/y$. Then
$\frac{\partial}{\partial x}\left(1-\frac{y}{x}\right)^{x}$ $=\frac{\partial z}{\partial x}\frac{\partial}{\partial z}\left(\left(1-\frac{1}{z}\right)^{z}\right)^{y}$ (By chain rule.)
$=\frac{y}{y}\left(\left(1-\frac{1}{z}\right)^{z}\right)^{y-1}\frac{\partial}{\partial z}\left(1-\frac{1}{z}\right)^{z}$ (By chain rule.)
$\geq 0\,.$ (Since $z>1$)
This concludes the proof of the lemma. ∎
Lemma 9 (Folklore).
Consider a reversible Markov chain on $G$ with transition matrix $\mathcal{P}$. For all $u,v\in V$ and $\ell\in\mathbbm{N}$, we have
$\mbox{\em E}_{visit}(u,v,\ell)\pi(u)=\mbox{\em E}_{visit}(v,u,\ell)\pi(v)\,.$
Proof.
We have
$\mbox{E}_{visit}(u,v,\ell)\pi(u)$ $=\sum_{i=1}^{\ell}\pi(u)\mathcal{P}^{i}_{u,v}$ (By definition of $\mbox{E}_{visit}$.)
$=\sum_{i=1}^{\ell}\pi(v)\mathcal{P}^{i}_{v,u}$ (By [LP17] equation 1.30.)
$=\mbox{E}_{visit}(v,u,\ell)\pi(v)\,.$ (By definition of $\mbox{E}_{visit}$.)
This concludes the proof of the lemma. ∎
We also use the following lemma from [BCR24].
Lemma 10 ([BCR24], Lemma 6).
Suppose $\mathbf{x}=(x_{0},x_{1},\ldots,x_{\ell})$ is a walk on $G$ and $f:V\to\mathbbm{R}$ is a valid function for the walk $\mathbf{x}$. Then $f$ has a unique local minimum at $x_{\ell}$, the last vertex on the walk.
Appendix B Corollaries of the main theorem
We can connect 1 to the spectral gap of the transition matrix of the Markov chain used via the following inequality  (see, e.g., [LP17], Theorem 12.4): If $\mathcal{P}$ is lazy, irreducible, and time-reversible, then
$t_{mix}(\epsilon)\leq\left(\frac{1}{1-\lambda_{2}}\right)\log\left(\frac{1}{\epsilon\min_{x\in V}\pi(x)}\right)\,.$ (171)
We obtain the following corrollary, which lower bounds the randomized complexity of local search as a function of the spectral gap of $\mathcal{P}$. See 1
Proof.
We proceed by substituting Eq. 171 into 1. This directly yields that the randomized query complexity of local search on $G$ is
$\Omega\left(\frac{(1-\lambda_{2})\sqrt{n}}{\log\big{(}2n/(\sigma\min_{x\in V}\pi(x))\big{)}\exp(3\sigma)}\right)\,.$ (172)
By definition of $\sigma$ we have
$\sigma\min_{x\in V}\pi(x)=\max_{x\in V}\pi(x)\geq 1/n\,.$ (173)
Combining Eq. 172 with Eq. 173 yields that the randomized query complexity of local search on $G$ is
$\Omega\left(\frac{(1-\lambda_{2})\sqrt{n}}{\log(n)\exp(3\sigma)}\right)\,.$ (174)
This completes the proof of the corollary. ∎
The prior work in [BCR24] implies a result similar to but weaker than 1. To get a result in terms of spectral gap via [BCR24], we need to connect the edge expansion $\beta$ to the spectral gap $1-\lambda_{2}$ of a particular Markov chain. We will do this via the bottleneck ratio $\Phi_{\star}$.
Lemma 11 ([LP17], Theorem 13.3).
If the Markov chain is lazy, then
$\frac{\Phi_{\star}^{2}}{2}\leq 1-\lambda_{2}\leq 2\Phi_{\star}\,.$ (175)
We can use this to get a bound on local search in terms of spectral gap via the following lemma from [BCR24].
Lemma 12 ([BCR24], corollary 2).
The randomized query complexity of local search is in
$\Omega\left(\frac{\beta\sqrt{n}}{d_{max}\log^{2}(n)}\right)$ (176)
As a special case, consider the simple lazy random walk on a graph with $d_{max}/d_{min}\leq C$ for constant $C$. Applying Lemma 12 to this walk yields the following result: See 1
Proof.
For the simple lazy random walk we have $\pi(u)=d(u)/(2|E|)$ for all $u\in V$. Let $\mathcal{P}$ be the transition matrix of the simple lazy random walk. Then
$\Phi_{\star}$ $=\min_{S\subseteq V\;\mid\;\pi(S)\leq 1/2}\frac{\sum_{(u,v)\in E(S,S^{c})}\pi(u)\mathcal{P}_{u,v}}{\pi(S)}$ (By definition of $\Phi_{\star}$.)
$=\min_{S\subseteq V\;\mid\;\pi(S)\leq 1/2}\frac{\sum_{(u,v)\in E(S,S^{c})}(d(u)/(2|E|))(1/d(u)))}{\sum_{u\in S}(d(u)/(2|E|))}$ (By definition of $\mathcal{P}$.)
$=\min_{S\subseteq V\;\mid\;\pi(S)\leq 1/2}\frac{|E(S,S^{c})|}{\sum_{u\in S}d(u)}\,.$ (177)
To get a term of $\beta$, we need to change the scope of $S$ from
$\{S\subseteq V\mid\pi(S)\leq 1/2\}$
to
$\{S\subseteq V\mid|S|\leq n/2\}\,.$
Let $S^{*}$ be a minimizing choice of $S$ from $\{S\subseteq V\mid|S|\leq n/2\}$ for $|E(S,S^{c})|/\sum_{u\in S}d(u)$. Then either $S^{*}$ or $S^{*c}$ (or both) will be in $\{S\subseteq V\mid\pi(S)\leq 1/2\}$. We analyze these two cases separately.
Case 1: $\pi(S^{*})\leq 1/2$.
Continuing from Eq. 177, this gives us
$\Phi_{\star}$ $\leq\frac{|E(S^{*},S^{*c})|}{\sum_{u\in S^{*}}d(u)}$ (Since $\pi(S^{*})\leq 1/2$.)
$=\min_{S\subseteq V\;\mid\;|S|\leq n/2}\frac{|E(S,S^{c})|}{\sum_{u\in S}d(u)}\,.$ (By definition of $S^{*}$.)
Case 2: $\pi(S^{*c})\leq 1/2$.
Continuing from Eq. 177, this gives us
$\Phi_{\star}$ $\leq\frac{|E(S^{*},S^{*c})|}{\sum_{u\in S^{*c}}d(u)}$ (Since $\pi(S^{*c})\leq 1/2$.)
$\leq\frac{|E(S^{*},S^{*c})|}{\sum_{u\in S^{*}}d(u)}\cdot C$ (Since $|S^{*}|\leq n/2\leq|S^{*c}|$ and by definition of $C$.)
$=\min_{S\subseteq V\;\mid\;|S|\leq n/2}\frac{|E(S,S^{c})|}{\sum_{u\in S}d(u)}\cdot C\,.$ (By definition of $S^{*}$.)
In both cases, we have
$\Phi_{\star}$ $\leq\min_{S\subseteq V\;\mid\;|S|\leq n/2}\frac{|E(S,S^{c})|}{\sum_{u\in S}d(u)}\cdot C$ (By LABEL:eq:phi_beta_case_1 and LABEL:eq:phi_beta_case_2)
$\leq\min_{S\subseteq V\;\mid\;|S|\leq n/2}\frac{|E(S,S^{c})|}{|S|\cdot d_{min}}\cdot C$ (By definition of $d_{min}$.)
$=\frac{\beta\cdot C}{d_{min}}$ (By definition of $\beta$.)
$\leq\frac{\beta\cdot C^{2}}{d_{max}}\,.$ (Since $d_{max}/d_{min}\leq C$.)
Substituting LABEL:eq:phi_beta into Lemma 11, we get
$1-\lambda_{2}\leq\frac{2\beta\cdot C^{2}}{d_{max}}\,.$ (181)
Substituting Eq. 181 into Lemma 12, we get a lower bound of $\Omega\bigl{(}\frac{(1-\lambda_{2})\sqrt{n}}{\log^{2}(n)}\bigr{)}$ on the randomized query complexity of local search on $G$. This completes the proof of the corollary. ∎ Compare to the following corollary, which is just 1 applied to the simple lazy random walk when $d_{max}/d_{min}$ is bounded by a constant $C$. See 2
Proof.
Since $d_{max}/d_{min}\leq C$, we have that $\exp(3\sigma)\leq\exp(3C)$, which is a constant. Therefore 1 directly gives that the randomized query complexity of local search is
$\Omega\left(\frac{(1-\lambda_{2})\sqrt{n}}{\log(n)\exp(3\sigma)}\right)=\Omega\left(\frac{(1-\lambda_{2})\sqrt{n}}{\log(n)}\right)\,.$ (182)
2 is stronger than 1 by a factor of $\log(n)$. This represents an improvement of this paper in bounding the difficulty of local search in terms of spectral gap.